OpenStudy (starryknight):

Solve for y in the following system of equations. -4x - 10y = 20 32x + 6y = -12 -2 -4 2 4 please help me ...... free metal for best answer

OpenStudy (texaschic101):

-4x - 10y = 20 -->(8) 32x + 6y = -12 ----------------- -32x - 80y = 160 (result of multiplying by 8) 32x + 6y = -12 -----------------add - 74y = 148 y = -148/74 y = - 2

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

x=0 y = 2

OpenStudy (texaschic101):

hmm..lets check...x = 0 and y = 2 -4x - 10y = 20 -4(0) - 10(2) = 20 -20 = 20 (no it does not) hazem....y is not 2, it is -2